# Markov Chain Variance Estimation: A Stochastic Approximation Approach

### v1

@@ Abstract
We consider the problem of estimating the asymptotic variance of a function defined on a Markov chain, an important step for statistical inference of the stationary mean. We design the first recursive estimator that requires $O(1)$ computation at each step, does not require storing any historical samples or any prior knowledge of run-length, and has optimal $O(\frac{1}{n})$ rate of convergence for the mean-squared error (MSE) with provable finite sample guarantees. Here, $n$ refers to the total number of samples generated. The previously best-known rate of convergence in MSE was $O(\frac{\log n}{n})$, achieved by jackknifed estimators, which also do not enjoy these other desirable properties. Our estimator is based on linear stochastic approximation of an equivalent formulation of the asymptotic variance in terms of the solution of the Poisson equation.
@@ 1 Introduction
Estimating $\kappa(f)$, while practically relevant, is particularly challenging compared to estimating variance using independent and identically-distributed (i.i.d.) samples because of temporal correlations introduced due to the underlying Markov chain structure. The problem of estimating $\kappa(f)$ has been extensively studied in literature. However, to the best of our knowledge, all the existing estimators have limitations of being either non-recursive in that the estimator at step $n$ cannot be expressed as a function of the various statistics computed at step $n-1$, necessitating re-computing the estimator from scratch at each step, or requiring to store a large number of historical observations, or having a sub-optimal rate of convergence of mean-squared estimation error to 00, or some combination of these. We discuss the related literature in details in Section 1.2. In this work, our goal is to design an estimator for $\kappa(\cdot)$ that utilizes samples from the given Markov chain along a single trajectory, is recursive and memory-efficient, easy to compute, and has an optimal rate of convergence in mean-squared sense.
@@ 1.1 Contributions
    We propose the first recursive estimator (or algorithm) for estimating $\kappa(f)$ that requires $O(1)$ computation at each step, does not require storing any historical samples, and enjoys optimal $O(\frac{1}{n})$ rate of convergence for the mean-squared error (MSE). Here, $n$ refers to the total number of samples generated. The best previously known rate-of-convergence in MSE for any estimator (irrespective of their memory and computation requirements or being recursive) was $O(\frac{\log n}{n})$ (cf., jackknifed estimators in Dingeç et al. (2015)).
@@ 1.2 Related Literature
We now discuss the existing estimators for $\kappa(f)$ and related literature in details.
Table 1: Summary of the results on estimators for $\kappa(f)$
Estimator Representative references (non exhaustive) Mean squared error Recursive[a] Memory requirement[b]
@@
Linear SA This work $O\left(\frac{1}{n}\right)$ $O(1)$
  • [a] The estimator at step $n$ can be expressed as a function of the various estimates at step $n-1$, and can be computed in a constant time (independent of $n$). A non recursive estimator requires re-computing the estimator from scratch at each step.
    [b] The memory required for computing the estimator at step $n$.
@@
A popular estimator for $\kappa(f)$ is the BM estimator Conway et al. (1959); Schmeiser (1982), which we informally describe next. Consider $n$ observations of the function $f$ evaluated on the $n$ consecutive states of the underlying Markov chain, and divide these observations into $b$ non-overlapping batches, each of size $m$ (for simplicity, assume that $n=mb$). For $i\in[b]$, let $m_{i}$ denote the sample mean of the observations from batch $i$. Then, the BM estimator corresponds to the sample variance of $m_{i}$ scaled by the batch size $m$. BM and its variants have been extensively studied (see, Schmeiser (1982); Meketon and Schmeiser (1984); Carlstein (1986); Welch (1987); Kunsch (1989); Alexopoulos et al. (2007); Meketon and Schmeiser (2007); Alexopoulos et al. (2011)). See Politis et al. (1999) for an introduction to the different estimators and their relations. Roughly speaking, here, the batch size $m$ governs the bias, while the number of batches $b$ governs the variance of the estimator.
@@
Several bootstrap-based estimators have also been proposed in literature Kunsch (1989); Liu et al. (1992); Carlstein (1986); Romano (1992); Politis and Romano (1994). However, these are computationally intensive, and hence, we do not discuss them in details in this work. Nordman (2009); Lahiri (1999) study the asymptotic rates for bias and variance of the different bootstrap-based estimators; results summarized in Table 1. We refer the reader to Lahiri (2013) for an introduction to this approach. Schruben (1983) develop standardized time series based estimators, extensively studied in Goldsman and Schruben (1984); Glynn and Iglehart (1988); Goldsman et al. (1990); Goldsman and Schruben (1990). Trevezas and Limnios (2009) propose maximum likelihood estimator for $\kappa(f)$, and prove that it is strongly consistent and asymptotically normal.
@@ 1.3 Setup and Asymptotic Variance
It is worth noting that the first term in (3) accounts for the per-step stationary variance, while the second term encompasses the temporal correlation introduced due to the Markov chain structure. These different representations for asymptotic variance are well known in literature. For example, see (Douc et al., 2018, Theorem 21.2.6) for a similar formulation. For completeness, we give a proof of the Proposition in Appendix B.1.
@@ 1.4 Stochastic Approximation: An Overview
As we will see in the following sections, these update methods play a crucial role in arriving at a simple recursive estimator that satisfies all the desirable properties listed earlier in Table 1.
@@ 2.2 Convergence Rates
Theorem 2.1$(a)$ bounds the mean-squared estimation error of the proposed algorithm in case of constant step size. Though the first term in this bound decays exponentially fast as the number of iterations $n$ increase, the other two terms are constant and become a bottleneck after sufficiently large $n$. This is a well-known behavior of SA algorithms with constant step sizes. In fact, this algorithm suffers similar drawback of being non-recursive as the vanilla Monte Carlo BM estimator (see, Table 1). In order to ensure the mean-squared error smaller than $\epsilon^{2}$ for $\epsilon>0$ (or mean absolute error smaller than $\epsilon$), one needs to pick the step size $\alpha$ as a function of $\epsilon$ (also see Corollary 2.1.1 below). This drawback of being non-recursive is overcome by choosing the step size to be diminishing (as in Theorem 2.1$(b)$).
@@
Recall that previously best-known rate of convergence for MSE for estimating $\kappa(f)$ for a Markov chain was $O(\frac{\log n}{n})$ (see Table 1). For diminishing step size, Algorithm 1 achieves ${O}(\frac{1}{n})$ rate of convergence (Theorem 2.1$(b)$). As in part $(a)$, the first term in the bound in part $(b)$ decays faster, and the other two terms are rate-determining $O(\frac{1}{n})$ terms. We believe that the rate of convergence of $\Omega(\frac{1}{n})$ for MSE is tight. In the special setting of i.i.d. noise with parametric distributions, this follows from the Cramér-Rao lower bound Nielsen (2013). We discuss a simple estimator that achieves this rate in i.i.d. setting in Appendix C.4, and also propose a SA-based recursive estimator achieving the same rate in i.i.d. setting in the following section (Section 2.3).
@@ 2.3 Estimating Stationary Variance
In the previous section, we saw an estimator for the asymptotic variance $\kappa(f)$, which captures the temporal correlations introduced due to the underlying Markov chain. Suppose you want to estimate the stationary variance, instead. Formally, this corresponds to estimating $v(f):=\mathbb{E}_{\pi\!}\!\left[(f(X)-\bar{f})^{2}\right].$ In Algorithm 2, we present a sequential (recursive) estimator for $v(f)$, that also satisfies all the properties listed in Table 1. Our estimator corresponds to SA update for the following equation:
@@
The corollary below shows that Algorithm 2 can also be used to estimate the variance in i.i.d. setting, achieving the optimal rate of convergence of $O(\frac{1}{n})$ in MSE. Moreover, the estimator satisfies all the desirable properties stated in Table 1 – it is recursive, does not require storing historical samples, and does not require knowing the total run length apriori.
@@ 3 Variance Estimation: Linear Function Approximation
When the underlying state and action spaces are large, estimating $V$ function for each state-action pair may be computationally intractable. To address this, we consider evaluating an approximation of $V$ by its projection on a linear subspace spanned by a given fixed set of $d$ vectors $\{\tilde{\phi}_{1},\dots,\tilde{\phi}_{d}\}$, where $\tilde{\phi}_{i}\in\mathbb{R}^{S}$ for each $i\in[d]$. More concretely, for $s\in{\mathcal{S}}$, and $\theta\in\mathbb{R}^{d}$, we consider the following linear function approximation (LFA) of $V(s)$:
@@ 6 Conclusions and Future Work
We proposed the first recursive estimator for the asymptotic variance of a Markov chain that requires $O(1)$ computation at each step, does not require storing historical samples, and enjoys the optimal $O(\frac{1}{n})$ rate of convergence for MSE. We use Stochastic Approximation to design the proposed estimator. Using Lyapunov drift arguments, and Poisson equation to handle the Markov noise in analysis, we established finite sample bounds on the performance of the proposed estimator. We generalized it in several directions, including covariance matrix estimation for vector-valued functions of a Markov chain and estimating the asymptotic variance of a given stationary policy in an average reward RL framework. Our tighter analysis using Poisson equations to handle the Markov noise improves the MSE bound for linear SA in presence of Markov noise, in general, and for policy evaluation problem in RL, in particular. We established the ${O}(\frac{1}{n})$ rate of convergence of the proposed algorithm in the tabular and linear function approximation settings. We also characterized the approximation error in the latter setting. Below, we discuss a few directions for future research.

### v2

@@ Abstract
We consider the problem of estimating the asymptotic variance of a function defined on a Markov chain, an important step for statistical inference of the stationary mean. We design a novel recursive estimator that requires $O(1)$ computation at each step, does not require storing any historical samples or any prior knowledge of run-length, and has optimal $O(\frac{1}{n})$ rate of convergence for the mean-squared error (MSE) with provable finite sample guarantees. Here, $n$ refers to the total number of samples generated. Our estimator is based on linear stochastic approximation of an equivalent formulation of the asymptotic variance in terms of the solution of the Poisson equation.
@@ 1 Introduction
Estimating $\kappa(f)$, while practically relevant, is particularly challenging compared to estimating variance using independent and identically-distributed (i.i.d.) samples because of temporal correlations introduced due to the underlying Markov chain structure. The problem of estimating $\kappa(f)$ has been extensively studied in literature. However, to the best of our knowledge, most of the existing estimators have limitations of being either non-recursive in that the estimator at step $n$ cannot be expressed as a function of the various statistics computed at step $n-1$, necessitating re-computing the estimator from scratch at each step, or requiring to store a large number of historical observations, or having a sub-optimal rate of convergence of mean-squared estimation error to 00, or lack finite sample guarantees, or some combination of these. We discuss the related literature in details in Section 1.2. In this work, our goal is to design an estimator for $\kappa(\cdot)$ that utilizes samples from the given Markov chain along a single trajectory, is recursive and memory-efficient, easy to compute, and has an optimal rate of convergence in mean-squared sense, with provable finite sample guarantees.
@@ 1.1 Contributions
    We propose a novel recursive estimator (or algorithm) for estimating $\kappa(f)$ that requires $O(1)$ computation at each step, does not require storing any historical samples, and enjoys optimal $O(\frac{1}{n})$ rate of convergence for the mean-squared error (MSE). Here, $n$ refers to the total number of samples generated.
@@ 1.2 Related Literature
We now discuss the existing estimators for $\kappa(f)$ in details, which have been extensively studied in literature. We remark that most of these previous estimators have been studied under more general setups than considered in this work. However, in this section, we present their properties for the specific setting of finite state DTMC. Furthermore, we point out that asymptotic variance has also been referred to as time average variance constant (TAVC) in several previous works.
Table 1: Summary of the results on estimators for $\kappa(f)^{[a]}$
Estimator Representative references (non exhaustive) MSE Recursive[b] Memory requirement[c]
@@
Regenerative Crane and Iglehart (1974) $O\left(\frac{1}{n}\right)^{[d]}$ $O(1)$
simulation Glynn and Iglehart (1987)
Linear SA This work $O\left(\frac{1}{n}\right)^{[e]}$ $O(1)$
  • [a]  Different works in literature estimate $\kappa(f)$ under different conditions on the underlying process. However, in this table, we present their rates and memory requirements for the setting of finite state DTMC.
    [b]  Estimator at step $n$ can be expressed as a function of various estimates at $n-1$, and can be computed in a constant time (independent of $n$). A non recursive estimator requires re-computing the estimator from scratch at each step.
    [c]  Memory required for computing the estimator at step $n$ (as a function of $n$).
    [d]  Asymptotic rates of convergence have been studied. However, finite-sample guarantees are not known.
    [e]  Finite sample bound on MSE.
@@
A popular estimator for $\kappa(f)$ is the BM estimator Conway et al. (1959); Schmeiser (1982), which we informally describe next. Consider $n$ observations of the function $f$ evaluated on the $n$ consecutive states of the underlying Markov chain, and divide these observations into $b$ non-overlapping batches, each of size $m$ (for simplicity, assume that $n=mb$). For $i\in(b)$, let $m_{i}$ denote the sample mean of the observations from batch $i$. Then, the BM estimator corresponds to the sample variance of $m_{i}$ scaled by the batch size $m$. BM and its variants have been extensively studied (see, Schmeiser (1982); Meketon and Schmeiser (1984); Carlstein (1986); Welch (1987); Kunsch (1989); Alexopoulos et al. (2007); Meketon and Schmeiser (2007); Alexopoulos et al. (2011)). See Politis et al. (1999) for an introduction to the different estimators and their relations. Roughly speaking, here, the batch size $m$ governs the bias, while the number of batches $b$ governs the variance of the estimator.
@@
Several bootstrap-based estimators have also been proposed in literature Kunsch (1989); Liu et al. (1992); Carlstein (1986); Romano (1992); Politis and Romano (1994). However, these are computationally intensive, and hence, we do not discuss them in details in this work. Nordman (2009); Lahiri (1999) study the asymptotic rates for bias and variance of the different bootstrap-based estimators. We refer the reader to Lahiri (2013) for an introduction to this approach. Schruben (1983) develop standardized time series based estimators, extensively studied in Goldsman and Schruben (1984); Glynn and Iglehart (1988); Goldsman et al. (1990); Goldsman and Schruben (1990); Foley and Goldsman (1999); Aktaran-Kalayci et al. (2011). Trevezas and Limnios (2009) propose maximum likelihood estimator for $\kappa(f)$, and prove that it is strongly consistent and asymptotically normal.
@@
Regenerative simulation
Another class of estimators that have been extensively studied at least since 1970s is the regeneration-based estimators. These depend on existence of regeneration points in the underlying process at which the process can be thought to probabilistically start afresh and evolve independent of the past. Note that in a finite state DTMC, the (random) times of successive visits of the Markov chain to a particular state can serve as the regeneration points. Hence, loosely speaking, a DTMC path can be thought of as several independent cycles pasted together. A number of past works have exploited this property of the underlying process to design efficient recursive estimators for $\kappa(f)$ (see, Crane and Iglehart (1974); Glynn and Iglehart (1987, 1993); Henderson and Glynn (2001); Glynn and Whitt (2002) and Glynn (2006) for a nice exposition), which can be implemented in an online fashion using $O(1)$ computation. Moreover, central limit theorems (CLTs) are established for these estimators, which establish an optimal asymptotic error rate of $O(\frac{1}{\sqrt{n}})$ (that is comparable to $O(\frac{1}{n})$ MSE). However, to the best of our knowledge, finite time bounds of these estimators have not been established. While the CLT strongly suggests that one should be able to obtain a finite time bound of $O(\frac{1}{n})$ MSE, the explicit dependence of the error on the parameters of the underlying chain is not known. In contrast, we propose an alternate approach to Markov chain variance estimation based on SA, and we present explicit finite time bounds on the MSE.
@@ 1.3 Setup and Asymptotic Variance
It is worth noting that the first term in (3) accounts for the per-step stationary variance, while the second term encompasses the temporal correlation introduced due to the Markov chain structure. These different representations for asymptotic variance are well known in literature at least since 1970s (Maigret, 1978). Also see (Meyn and Tweedie, 2012, Theorem 17.5.3), (Douc et al., 2018, Theorem 21.2.6), and more recently, Glynn and Infanger (2024), for similar formulations. For completeness, we give a proof of the Proposition in Appendix B.1.
@@ 1.4 Stochastic Approximation: An Overview
As we will see in the following sections, these update methods play a crucial role in arriving at a simple recursive estimator that satisfies all the desirable properties discussed earlier in Section 1.
@@ 2.2 Convergence Rates
Theorem 2.1$(a)$ bounds the mean-squared estimation error of the proposed algorithm in case of constant step size. Though the first term in this bound decays exponentially fast as the number of iterations $n$ increase, the other two terms are constant and become a bottleneck after sufficiently large $n$. This is a well-known behavior of SA algorithms with constant step sizes. In fact, this algorithm suffers similar drawback of being non-recursive as the vanilla Monte Carlo BM estimator. In order to ensure the mean-squared error smaller than $\epsilon^{2}$ for $\epsilon>0$ (or mean absolute error smaller than $\epsilon$), one needs to pick the step size $\alpha$ as a function of $\epsilon$ (also see Corollary 2.1.1 below). This drawback of being non-recursive is overcome by choosing the step size to be diminishing (as in Theorem 2.1$(b)$).
@@
For diminishing step size, Algorithm 1 achieves ${O}(\frac{1}{n})$ rate of convergence (Theorem 2.1$(b)$). As in part $(a)$, the first term in the bound in part $(b)$ decays faster, and the other two terms are rate-determining $O(\frac{1}{n})$ terms. We believe that the rate of convergence of $\Omega(\frac{1}{n})$ for MSE is tight. In the special setting of i.i.d. noise with parametric distributions, this follows from the Cramér-Rao lower bound Nielsen (2013). We discuss a simple estimator that achieves this rate in i.i.d. setting in Appendix C.4, and also propose a SA-based recursive estimator achieving the same rate in i.i.d. setting in the following section (Section 2.3).
@@ 2.3 Estimating Stationary Variance
In the previous section, we saw an estimator for the asymptotic variance $\kappa(f)$, which captures the temporal correlations introduced due to the underlying Markov chain. Suppose you want to estimate the stationary variance, instead. Formally, this corresponds to estimating $v(f):=\mathbb{E}_{\pi\!}\!\left[(f(X)-\bar{f})^{2}\right].$ In Algorithm 2, we present a sequential (recursive) estimator for $v(f)$, that also satisfies all the properties discussed in Section 1. Our estimator corresponds to SA update for the following equation:
@@
The corollary below shows that Algorithm 2 can also be used to estimate the variance in i.i.d. setting, achieving the optimal rate of convergence of $O(\frac{1}{n})$ in MSE. Moreover, the estimator satisfies all the desirable properties discussed in Section 1 – it is recursive, does not require storing historical samples, and does not require knowing the total run length apriori.
@@ 3 Variance Estimation: Linear Function Approximation
When the underlying state and action spaces are large, estimating $V$ function for each state-action pair may be computationally intractable. To address this, we consider evaluating an approximation of $V$ by its projection on a linear subspace spanned by a given fixed set of $d$ vectors $\{\tilde{\phi}_{1},\dots,\tilde{\phi}_{d}\}$, where $\tilde{\phi}_{i}\in\mathbb{R}^{S}$ for each $i\in(d)$. More concretely, for $s\in{\mathcal{S}}$, and $\theta\in\mathbb{R}^{d}$, we consider the following linear function approximation (LFA) of $V(s)$:
@@ 6 Conclusions and Future Work
We proposed a novel recursive estimator for the asymptotic variance of a Markov chain that requires $O(1)$ computation at each step, does not require storing historical samples, and enjoys the optimal $O(\frac{1}{n})$ rate of convergence for MSE. We use Stochastic Approximation to design the proposed estimator. Using Lyapunov drift arguments, and Poisson equation to handle the Markov noise in analysis, we established finite sample bounds on the performance of the proposed estimator. We generalized it in several directions, including covariance matrix estimation for vector-valued functions of a Markov chain and estimating the asymptotic variance of a given stationary policy in an average reward RL framework. Our tighter analysis using Poisson equations to handle the Markov noise improves the MSE bound for linear SA in presence of Markov noise, in general, and for policy evaluation problem in RL, in particular. We established the ${O}(\frac{1}{n})$ rate of convergence of the proposed algorithm in the tabular and linear function approximation settings. We also characterized the approximation error in the latter setting. Below, we discuss a few directions for future research.
@@
  • In Section 1.2, we mentioned that the asymptotic behavior of regenerative simulation based estimators for $\kappa(f)$ has been extensively studied in literature. CLTs for these estimators have been proven, which can be used to establish an optimal asymptotic rate of $O(\frac{1}{n})$ in MSE. Investigating the finite sample behavior of these estimators and comparing its performance to our SA-based estimator is an interesting direction for future research.